OpenStudy (anonymous):

Solve by the elimination method. 2x + 3y = 18 5x – y = 11

OpenStudy (anonymous):

Give me a second.

OpenStudy (anonymous):

okey dokey :)

OpenStudy (anonymous):

5(2x + 3y) = 18 and 2(5x - y) = 11 Do you understand why I did that?

OpenStudy (anonymous):

So now you are left with 10x + 15y = 18 and 10x - 2y = 11

OpenStudy (anonymous):

The two "x"s cancel out so you have 17y = 7

OpenStudy (anonymous):

divide and what does x equal

OpenStudy (anonymous):

but now you need to find Y

OpenStudy (anonymous):

so substitute what you got for X in to find Y

OpenStudy (anonymous):

oh ok i kind of understand 7/17?

OpenStudy (anonymous):

for x?